# Schatten properties of commutators of fractional integrals on spaces of homogeneous type

### v1

@@ 1. Introduction
see [16, Chapter 6]. We write the two equivalent formulas above, parametrised by $\alpha$ and $\varepsilon$, since they give rise to two different classed of generalisations, as we will see below.
@@
where $a_{n}(R)$ is the $n$th approximation number (or singular value) of $R:L^{2}(\mu)\to L^{2}(\mu)$. In this direction, Janson and Peetre [26, p. 484] obtained the following results in the Euclidean space $\mathbb{R}^{d}$ with $d\geq 2$ as a special case of their work on so-called “paracommutators”:
@@
Many works on fractional integrals over spaces of homogeneous type, like [14], are formulated in so-called “normal” spaces with $V(x,y)\sim\rho(x,y)$, in which case the volumic and the metric versions coincide. Without assuming normality, volumic fractional integrals have been studied e.g. in [31], and [5] obtained the $(L^{p},L^{q})$ boundedness of their commutators $b,I_{\varepsilon}]$ for $b\in BMO(X)$. These volumic fractional integrals admit a relatively clean theory in its own right. Nevertheless, it seems that metric fractional integrals, especially with $\alpha=1$, are actually the ones that more frequently arise in applications; see e.g. [19, Theorem 3.22 and Section 9.1] and [32, Eq. (7)], where further references to such operators in different contexts are given. In particular, fractional powers $(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ of the Bessel Laplacian $\Delta_{\lambda}$ turn out to be of the metric form; see [3] and Section 12 below. Incorporating this prominent example into our theory was a major motivation for dealing with $\tilde{I}_{\alpha}$, and we will return to this example in more detail below. More generally, we show in Section 11 that a large class of fractional operators arising from heat kernels fall under the umbrella of metric fractional integrals.
@@
    $(\mathbb{R}^{n+1}_{+},|\cdot|,dm_{\lambda}^{(n+1)})$ is a space of homogeneous type of lower dimension $d=n+1>2$ that satisfies the $(1,2)$ (in fact, even the stronger $(1,1)$) Poincaré inequality. This is [23, Proposition 4.2], restated as Proposition 12.3.
@@ 2.1. Fractional integrals and their commutators
Although “strongly non-degenerate” is the main notion of non-degeneracy that we use in this paper, we reserve the simpler name “non-degenerate” for another variant (Definition 10.1), since this variant is closer to the notion of “non-degenerate” for singular integral kernels as defined in [23, Eq. (1.7)].
The existence of a (strongly) non-degenerate kernel requires in particular that $B(x,r)\neq X$ for all $r>0$, hence that $X$ is unbounded, and thus (by [5, Lemma 1.9]) that $\mu(X)=\infty$. Hence, this assumption is implicitly in force in all results dealing with (strongly) non-degenerate kernels. A modification can be made to accommodate spaces of finite diameter; see [23, Section 10] for related discussion in the case of singular (instead of fractional) kernels.
Moreover, the existence of a (strongly) non-degenerate kernel also implies the non-empty annulus property $B(x,\overline{C}r)\setminus B(x,r)\neq\varnothing$ for all $x\in X$ and $r>0$, which is equivalent to inequality (2.4) for some $d>0$ by [18, Remark 1.2]. In particular, this implies that $\mu(\{x\})=0$ for all $x\in X$, which is hence also implicitly assumed in all results dealing with (strongly) non-degenerate kernels.
@@ 4. The upper bounds for the fractional commutators
For $p\in(2,\infty)$, this is [23, Proposition 5.8]. For $p=2$, an explicit formula of the Hilbert-Schmidt norm of integral operators (see e.g. [34, Theorem 2.11]) and (4.1) give that
@@ 5. Characterization for fractional oscillation spaces and Besov spaces
For the fractional version, an analogous is stated below in Propositions 5.2. First, we give a useful characterisation from [23, Lemma 11.4]:
@@
Taking $r=p$ and $s=1$ in [23, Proposition 11.3], we get for any given $Q\in\mathscr{D}$,
@@ 6. The complex median method of Wei–Zhang revisited
Two methods for proving lower bounds for commutators in the recent literature are the “approximate weak factorisation” from [22], and versions of the so-called median method. Until recently, this second approach was restricted to real-valued functions, for which the median is conventionally defined. However, Wei and Zhang [36] recently demonstrated that, with suitable modifications, the median method can also be extended to work with complex-valued functions. Once their [36, Theorem 1.5] on the existence of the complex median is proved, it can be conveniently applied as a black box; however, the proof of [36, Theorem 1.5] itself is not so easy. For this reason, we present below a variant of this theorem, with a much simpler proof and roughly the same scope of applicability.
@@
Recall that the median of a probability $\nu$ on $\mathbb{R}$ is a value $m\in\mathbb{R}$ such that both $\nu((-\infty,m])\geq\frac{1}{2}$ and $\nu([m,\infty))\geq\frac{1}{2}$, i.e., the mass of $\nu$ essentially equally divided on two sides of $m$, to the extent allowed by a possible point mass at $m$.
The value $m\in\mathbb{C}$ in Proposition 6.1 is a “quasi-median” of $\nu$ in the sense that the measure of $\nu$ is divided among each of the $N$ conical regions around $m$, with a fair (if not quite equal) share to each.
Proposition 6.1 also holds in each of the cases $(N,\varepsilon)\in\{(1,1),(2,\tfrac{1}{2}),(4,\tfrac{1}{16})\}$, where case $(N,\varepsilon)=(1,1)$ is the triviality $\nu(\mathbb{C})\geq 1$, while case $N=2$ and $\varepsilon=\frac{1}{2}$ is a simple extension of the existence of the median for the probability $\nu_{1}(A):=\nu(A\times\mathbb{R})$ on $\mathbb{R}$: if $m$ is a median of $\nu_{1}$, the required property is satisfied with the cones $\Gamma(\pm 1,\pi)$.
The nontrivial case $(N,\varepsilon)=(4,\frac{1}{16})$ is a recent result of Wei and Zhang, [36, Theorem 1.5]. Its proof is “elementary” but rather tedious, taking about 10 pages in [36, Section 6].
We will show that case $(N,\varepsilon)=(3,\frac{1}{4})$ holds with a much simpler proof, while the result has roughly the same scope of applicability as [36, Theorem 1.5] for commutator estimates, as we will see further below.
The proof below shows that the cones can be chosen to be in the standard orientation $\Gamma_{j}=\Gamma(e^{ij\frac{2\pi}{3}},\frac{2\pi}{3})$ (or, by a rotation, in any other prescribed orientation that we like). This is in contrast to [36, Theorem 1.5], where the orientation of the cones may depend on $\nu$.
@@ 7. The lower bounds for the fractional commutators
We will also need the following result, which is [23, Corollary 7.7]:
@@ 8. Characterisation of constants
Motivated by the result from [25, Proposition 4.1], we give a characterisation of constants via Besov-type conditions, with the assumption of the Poincaré inequality.
@@ 10. Fractional integrals with regular kernels
Our main results in this paper, whose proofs we have just completed, only depend on the size and not on the any regularity of the kernel $K$ of the fractional integral. Nevertheless, many examples of fractional integrals arising in applications also possess additional regularity, and we hence some comments on this situation. In particular, we observe that, with some mild regularity of the kernel, the strong non-degeneracy that we have assumed (a uniform estimate over all points in certain balls) already follows from a simpler non-degeneracy condition involving a pair of points only.
@@ 12. Application to fractional Bessel operators
In [3], Betancor et al. gave the kernel estimates of the one-dimensional fractional Bessel operator $(-\Delta_{\lambda}^{(1)})^{-\alpha/2}$ on $(0,\infty)$ with parameters $n=0$, $\lambda>0$ and $0<\alpha<1+2\lambda$. In this section, we extend their result to the corresponding fractional Bessel operator $(-\Delta_{\lambda}^{(n+1)})^{-\alpha/2}$ for all $n\geq 0$, $\lambda>0$ and $0<\alpha<n+1+2\lambda$. Taking $\mathcal{L}=-\Delta_{\lambda}$ and $s=\alpha/2$ in (11.1), the fractional Bessel operator is defined by
@@
It remains to show the relevant conditions for the kernel of the fractional Bessel operators. We will derive these from bounds for the heat kernels associated to Bessel operators from [10, Section 7.5].
@@
As for the case $n=0$, the size estimate and the smoothness estimate of the kernel $K_{{\lambda,\alpha}}$ associated with the fractional Bessel operator have been proved by Betancor et.al. in [3, Lemmas 5.1, 5.2]. The following Proposition 12.5 extends their result to all $n\geq 0$ and gives the lower bound of the kernel $K_{{\lambda,\alpha}}$.

### v2

@@ 1. Introduction
see [17, Chapter 6]. We write the two equivalent formulas above, parametrised by $\alpha$ and $\varepsilon$, since they give rise to two different classes of generalisations, as we will see below.
@@
where $a_{n}(R)$ is the $n$th approximation number (or singular value) of $R:L^{2}(\mu)\to L^{2}(\mu)$. In this direction, Janson and Peetre [27, p. 484] obtained the following results in the Euclidean space $\mathbb{R}^{d}$ with $d\geq 2$ as a special case of their work on so-called “paracommutators”:
@@
Many works on fractional integrals over spaces of homogeneous type, like [15], are formulated in so-called “normal” spaces with $V(x,y)\sim\rho(x,y)$, in which case the volumic and the metric versions coincide. Without assuming normality, volumic fractional integrals have been studied e.g. in [32], and [6] obtained the $(L^{p},L^{q})$ boundedness of their commutators $b,I_{\varepsilon}]$ for $b\in BMO(X)$. These volumic fractional integrals admit a relatively clean theory in its own right. Nevertheless, it seems that metric fractional integrals, especially with $\alpha=1$, are actually the ones that more frequently arise in applications; see e.g. [20, Theorem 3.22 and Section 9.1] and [33, Eq. (7)], where further references to such operators in different contexts are given. In particular, fractional powers $(-\Delta_{\lambda})^{-\frac{\alpha}{2}}$ of the Bessel Laplacian $\Delta_{\lambda}$ turn out to be of the metric form; see [4] and Section 12 below. Incorporating this prominent example into our theory was a major motivation for dealing with $\tilde{I}_{\alpha}$, and we will return to this example in more detail below. More generally, we show in Section 11 that a large class of fractional operators arising from heat kernels fall under the umbrella of metric fractional integrals.
@@
    $(\mathbb{R}^{n+1}_{+},|\cdot|,dm_{\lambda}^{(n+1)})$ is a space of homogeneous type of lower dimension $d=n+1>2$ that satisfies the $(1,2)$ (in fact, even the stronger $(1,1)$) Poincaré inequality. This is [24, Proposition 4.2], restated as Proposition 12.3.
@@ 2.1. Fractional integrals and their commutators
Although “strongly non-degenerate” is the main notion of non-degeneracy that we use in this paper, we reserve the simpler name “non-degenerate” for another variant (Definition 10.1), since this variant is closer to the notion of “non-degenerate” for singular integral kernels as defined in [24, Eq. (1.7)].
The existence of a (strongly) non-degenerate kernel requires in particular that $B(x,r)\neq X$ for all $r>0$, hence that $X$ is unbounded, and thus (by [6, Lemma 1.9]) that $\mu(X)=\infty$. Hence, this assumption is implicitly in force in all results dealing with (strongly) non-degenerate kernels. A modification can be made to accommodate spaces of finite diameter; see [24, Section 10] for related discussion in the case of singular (instead of fractional) kernels.
Moreover, the existence of a (strongly) non-degenerate kernel also implies the non-empty annulus property $B(x,\overline{C}r)\setminus B(x,r)\neq\varnothing$ for all $x\in X$ and $r>0$, which is equivalent to inequality (2.4) for some $d>0$ by [19, Remark 1.2]. In particular, this implies that $\mu(\{x\})=0$ for all $x\in X$, which is hence also implicitly assumed in all results dealing with (strongly) non-degenerate kernels.
@@ 4. The upper bounds for the fractional commutators
For $p\in(2,\infty)$, this is [24, Proposition 5.8]. For $p=2$, an explicit formula of the Hilbert-Schmidt norm of integral operators (see e.g. [35, Theorem 2.11]) and (4.1) give that
@@ 5. Characterization for fractional oscillation spaces and Besov spaces
For the fractional version, an analogous is stated below in Propositions 5.2. First, we give a useful characterisation from [24, Lemma 11.4]:
@@
Taking $r=p$ and $s=1$ in [24, Proposition 11.3], we get for any given $Q\in\mathscr{D}$,
@@ 6. The complex median method of Wei–Zhang revisited
Two methods for proving lower bounds for commutators in the recent literature are the “approximate weak factorisation” from [23], and versions of the so-called median method. Until recently, this second approach was restricted to real-valued functions, for which the median is conventionally defined. However, Wei and Zhang [37] recently demonstrated that, with suitable modifications, the median method can also be extended to work with complex-valued functions. A key to that extension is the existence of a “two-dimensional median” in a suitable sense. This existence is recently due to Baringhaus and Grübel [1], and a slightly weaker variant (but sufficient for commutators) was independently rediscovered by [37, Theorem 1.5]. (See Remark 6.2 below for details.) In Proposition 6.1 below, we present yet another variant of the existence of a two-dimensional median, with a simpler proof and roughly the same scope of applicability as far as commutator applications are concerned.
@@
Recall that the median of a probability $\nu$ on $\mathbb{R}$ is a value $m\in\mathbb{R}$ such that both $\nu((-\infty,m])\geq\frac{1}{2}$ and $\nu([m,\infty))\geq\frac{1}{2}$, i.e., the mass of $\nu$ is essentially equally divided on two sides of $m$, to the extent allowed by a possible point mass at $m$.
The value $m\in\mathbb{C}$ in Proposition 6.1 is a “quasi-median” of $\nu$ in the sense that the measure of $\nu$ is divided among each of the $N$ conical regions around $m$, with a fair (if not quite equal) share to each. We do not know whether Proposition 6.1 holds with $(N,\varepsilon)=(3,\frac{1}{3})$. This would be interesting in itself, but it makes no difference to our applications to commutators further below.
Proposition 6.1 also holds in each of the cases $(N,\varepsilon)\in\{(1,1),(2,\tfrac{1}{2}),(4,\tfrac{1}{4})\}$, where case $(N,\varepsilon)=(1,1)$ is the triviality $\nu(\mathbb{C})\geq 1$, while case $N=2$ and $\varepsilon=\frac{1}{2}$ is a simple extension of the existence of the median for the probability $\nu_{1}(A):=\nu(A\times\mathbb{R})$ on $\mathbb{R}$: if $m$ is a median of $\nu_{1}$, the required property is satisfied with the cones $\Gamma(\pm 1,\pi)$.
The nontrivial case $(N,\varepsilon)=(4,\frac{1}{4})$ is a recent result of Baringhaus and Grübel [1, Theorem 1], while Wei and Zhang [37, Theorem 1.5] independently obtain the weaker variant with $(N,\varepsilon)=(4,\frac{1}{16})$. The proof of [37, Theorem 1.5] is “elementary” but rather tedious, taking about 10 pages in [37, Section 6]. The proof of [1, Theorem 1] in [1, Section 4.1] takes less than a page, but it uses as input the special case of absolutely continuous measures, which is quoted from elsewhere.
We will show that case $(N,\varepsilon)=(3,\frac{1}{4})$ holds with a much simpler proof, while the result has roughly the same scope of applicability as [37, Theorem 1.5] for commutator estimates, as we will see further below.
The proof below shows that the cones can be chosen to be in the standard orientation $\Gamma_{j}=\Gamma(e^{ij\frac{2\pi}{3}},\frac{2\pi}{3})$ (or, by a rotation, in any other prescribed orientation that we like). This is in contrast to [1, Theorem 1] and [37, Theorem 1.5], where the orientation of the cones will also depend on $\nu$.
@@ 7. The lower bounds for the fractional commutators
We will also need the following result, which is [24, Corollary 7.7]:
@@ 8. Characterisation of constants
Motivated by the result from [26, Proposition 4.1], we give a characterisation of constants via Besov-type conditions, with the assumption of the Poincaré inequality.
@@ 10. Fractional integrals with regular kernels
Our main results in this paper, whose proofs we have just completed, only depend on the size and not on the any regularity of the kernel $K$ of the fractional integral. Nevertheless, many examples of fractional integrals arising in applications also possess additional regularity, and we hence make some comments on this situation. In particular, we observe that, with some mild regularity of the kernel, the strong non-degeneracy that we have assumed (a uniform estimate over all points in certain balls) already follows from a simpler non-degeneracy condition involving a pair of points only.
@@ 12. Application to fractional Bessel operators
In [4], Betancor et al. gave the kernel estimates of the one-dimensional fractional Bessel operator $(-\Delta_{\lambda}^{(1)})^{-\alpha/2}$ on $(0,\infty)$ with parameters $n=0$, $\lambda>0$ and $0<\alpha<1+2\lambda$. In this section, we extend their result to the corresponding fractional Bessel operator $(-\Delta_{\lambda}^{(n+1)})^{-\alpha/2}$ for all $n\geq 0$, $\lambda>0$ and $0<\alpha<n+1+2\lambda$. Taking $\mathcal{L}=-\Delta_{\lambda}$ and $s=\alpha/2$ in (11.1), the fractional Bessel operator is given by
@@
It remains to show the relevant conditions for the kernel of the fractional Bessel operators. We will derive these from bounds for the heat kernels associated to Bessel operators from [11, Section 7.5].
@@
As for the case $n=0$, the size estimate and the smoothness estimate of the kernel $K_{{\lambda,\alpha}}$ associated with the fractional Bessel operator have been proved by Betancor et.al. in [4, Lemmas 5.1, 5.2]. The following Proposition 12.5 extends their result to all $n\geq 0$ and gives the lower bound of the kernel $K_{{\lambda,\alpha}}$.